You and your friend, Jake, go to the roof of a 160 foot tall building. You drop your football over the edge at the same instant that Jake throws his football straight down at 48 feet per second. By how many seconds does his football beat yours to the ground? Using h=-16t^2+vt+c, write and solve tow equations to answer the question.
Answer by nerdybill(7384)

You and your friend, Jake, go to the roof of a 160 foot tall building. You drop your football over the edge at the same instant that Jake throws his football straight down at 48 feet per second. By how many seconds does his football beat yours to the ground? Using h=-16t^2+vt+c, write and solve tow equations to answer the question.
.
looking at the given formula:
h=-16t^2+vt+c
h is the height at time t (set to 0 for the ground)
v is the initial velocity or speed
c is the initial height
.
Jake:
0=-16t^2+48t+160
0=t^2-3t-10
0=(t-5)(t+2)
t = {-2, 5}
throw out the negative solution leaving
t = 5 secs
.
You:
0=-16t^2+0t+160
0=-16t^2+160
0=t^2-10
t^2 = 10
t = sqrt(10) secs
.
Time Jake beats you:
5-sqrt(10) = 1.84 seconds